OpenStudy (anonymous):

Finding x and y intercepts. Please help me understand this for my test that I have tomorrow. 3y-x=6 or 5x-6y=30

OpenStudy (anonymous):

Okay do you know when x and y intercepts occur?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

wait nvm i think i do kinda.

OpenStudy (anonymous):

but no whole lot tho

OpenStudy (anonymous):

An x intercept occur when the line crosses the x axis or y=0. A y intercept occurs when the line crosses the y axis or x=0

OpenStudy (anonymous):

Let's take 5x-6y=30 To find x-intercept plug y = 0 in the given equation. So, 5x-6(0)=30 5x=30 x = 6 is the x-intercept. Similarly plug x = 0 to find the y-intercept 5(0)-6y=30 -6y=30 y = -5 is the y -intercept... Hope this will help you

OpenStudy (anonymous):

This said we turn back to your original equation and plug in the value of zero in for x and solve for y. Then you do the reverse plug in zero for y and solve for x.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are always welcome :)